# Estimators for multivariate allometric regression model

### v2

@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=o(r_{n}^{2})$
for some non-decreasing sequence $\{r_{n}\}_{n\in\mathbb{N}}$ of positive values $0<r_{1}\leq r_{2}\leq\cdots$ satisfying that $r_{n}(\hat{w}^{*}-w^{*})$ is uniformly tight, the same conclusion follows.
@@
Suppose that, as $n\to\infty$, $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix with other parameters held fixed. Then, as $n\to\infty$, (6.1) holds.
@@ 7 Numerical simulations
First, we consider a traditional large sample regime and a weak identifiability regime, as follows. The number of response variables $p$ is 10, the number of explanatory variables $q$ is 5, and the sample size $n$ is $20,50,100,200,500$. The explanatory variables matrix $\bm{X}$ ($n\times q$ matrix) is a random sample of size $n$ drawn from $\mathcal{N}_{q}(\bm{0}_{q},\bm{I}_{q})$ with centralization ($\bm{X}^{\prime}\bm{1}_{n}=\bm{0}_{q}$). Let $\bm{\alpha}=\bm{1}_{q}$ (the $q\times 1$ vector whose all elements are $1$). Let $\bm{\Sigma}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}$, where we set the diagonal matrix of eigenvalues $\bm{\Lambda}={\rm diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ by $\lambda_{1}=2$ (traditional large sample regime) or $\lambda_{1}=1+1/n$ (weak identifiability regime) and by $\lambda_{i}=1,\ i=2,3,\dots,p$ and we also set the orthogonal matrix of eigenvectors $\bm{\Gamma}=(\bm{\gamma}_{1},\bm{\gamma}_{2},\dots,\bm{\gamma}_{p})$ by using normalized eigenvectors of a sample covariance matrix of the random sample of size $2p$ drawn from $\mathcal{N}_{p}(\bm{0}_{p},\bm{I}_{p})$. We note that the setting of $\bm{\Gamma}$ does not affect the results of simulations. Table 1-(a) (traditional large sample regime) and (b) (weak identifiability regime) show the results of mean squared errors of the estimators based on 1000 numerical simulations.
@@
We see from Table 1-(a) that $\hat{\bm{\gamma}}_{1}(0.3)$ is the best for all cases and that $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are better than both $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}^{(T)}$. These results suggest that the optimal value of $w$, which is unknown in practice, is in $(0,1/2)$ and motivate to use the proposed estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$. As previously anticipated, $\hat{\bm{\gamma}}_{1}^{(E)}$ is not better than other estimators.
@@
We see from Table 1-(b) that $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are the best for almost all cases. As an exception, for $n=20$, $\hat{\bm{\gamma}}_{1}(0.1)$ is the best, and $\hat{\bm{\gamma}}_{1}(w^{*})$ is slightly better than $\hat{\bm{\gamma}}_{1}^{(R)}$. As for $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, it is not good as $\hat{\bm{\gamma}}_{1}(w^{*})$ for small $n$, but their difference goes to 00 as $n$ increases. In contrast, $\hat{\bm{\gamma}}_{1}^{(E)}$ is quite bad, which may be due to the fact from Theorem 4.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ is not necessarily consistent because $\eta=1$ does not satisfy $\eta<1/2$. The inconsistency of $\hat{\bm{\gamma}}_{1}^{(E)}$ may affect the performance of $\hat{\bm{\gamma}}_{1}^{(T)}$, which is worse than $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, and $\hat{\bm{\gamma}}_{1}(w^{*})$.
@@
We see from Table 1-(c) that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the best for all cases. Although $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are not the best, their performances become closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase. As for $\hat{\bm{\gamma}}_{1}^{(E)}$, its performance is bad for small $p$ and rapidly closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase, which may be due to the fact from Theorem 5.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ is consistent because $\delta=0.8$ and $\beta=0.8$ satisfy $\delta>2-2\beta$. In contrast, $\hat{\bm{\gamma}}_{1}^{(R)}$ is bad especially for larger $p$, which may be due to the fact from Theorem 5 that $\hat{\bm{\gamma}}_{1}^{(R)}$ is not necessarily consistent because $\delta=0.8$ does not satisfy $\delta>1$.
In our settings, $\hat{\bm{\gamma}}_{1}^{(T)}$ and $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ are good in all cases examined. In particular, on the one hand $\hat{\bm{\gamma}}_{1}^{(T)}$ is better in typical settings; on the other hand $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is better in settings where $\lambda_{1}-\lambda_{2}$ is small.

### v4

@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
$\frac{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+a(n-1)+c\{{\rm tr}(\bm{\Sigma})-(n-2q)\lambda_{1}\}}{\left[d\{q+(n-1-2q)w^{*}\}+c\right]^{2}}=o(r_{n}^{2})$
for some non-decreasing sequence $\{r_{n}\}_{n\in\mathbb{N}}$ of positive values $0<r_{1}\leq r_{2}\leq\cdots$ satisfying that $r_{n}(\hat{w}^{*}-w^{*})$ is uniformly tight, the same conclusion follows. Here, the notation (3.2) is used.
@@
Suppose that $\bm{\alpha}\neq\bm{0}_{q}$ and that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix as $n\to\infty$ with other parameters held fixed. Then, as $n\to\infty$, (6.1) holds.
@@ 7 Numerical simulations
First, we consider a traditional large sample regime and a weak identifiability regime, as follows. The number of response variables $p$ is 10, the number of explanatory variables $q$ is 5, and the sample size $n$ is $20,50,100,200,500$. The explanatory variables matrix $\bm{X}$ ($n\times q$ matrix) is a random sample of size $n$ drawn from $\mathcal{N}_{q}(\bm{0}_{q},\bm{I}_{q})$ with centralization ($\bm{X}^{\prime}\bm{1}_{n}=\bm{0}_{q}$). Let $\bm{\alpha}=\bm{1}_{q}$ (the $q\times 1$ vector whose all elements are $1$). Let $\bm{\Sigma}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}$, where we set the diagonal matrix of eigenvalues $\bm{\Lambda}={\rm diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ by $\lambda_{1}=2$ (traditional large sample regime) or $\lambda_{1}=1+n^{-\eta}$ with $\eta=1/3,1/2,1$ (weak identifiability regime) and by $\lambda_{i}=1,\ i=2,3,\dots,p$ and we also set the orthogonal matrix of eigenvectors $\bm{\Gamma}=(\bm{\gamma}_{1},\bm{\gamma}_{2},\dots,\bm{\gamma}_{p})$ by using normalized eigenvectors of a sample covariance matrix of the random sample of size $2p$ drawn from $\mathcal{N}_{p}(\bm{0}_{p},\bm{I}_{p})$. We note that the setting of $\bm{\Gamma}$ does not affect the results of simulations. Table 1 (traditional large sample regime) and Table 2 (weak identifiability regime) show the results of mean squared errors of the estimators based on 1000 numerical simulations.
Table 1: The values of the mean squared errors (with the average of $\hat{w}^{*}$ and $w^{*}$) in the case under a traditional large sample regime.
$n$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.10517 0.03734 0.01831 0.00862 0.00349
$\hat{\bm{\gamma}}_{1}^{(E)}$ 0.90481 0.46512 0.22117 0.09505 0.03515
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.10723 0.03887 0.01873 0.00886 0.00363
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.10404 0.03740 0.01804 0.00852 0.00349
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.10113 0.03613 0.01745 0.00823 0.00337
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.09915 0.03531 0.01711 0.00806 0.00330
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.09949 0.03540 0.01724 0.00812 0.00331
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.12335 0.04309 0.02127 0.01000 0.00400
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.10101 0.03684 0.01782 0.00843 0.00344
$\hat{w}^{*}$ (0.19359) (0.12519) (0.11918) (0.12590) (0.13702)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.10109 0.03642 0.01762 0.00832 0.00341
$w^{*}$ (0.19237) (0.17417) (0.16823) (0.16535) (0.16363)
The best result up to significant digits in each setting is in bold face.
Table 2: The values of the mean squared errors (with the average of $\hat{w}^{*}$ and $w^{*}$) in the cases under weak identifiability regimes.
(a) Results when $\lambda_{1}-\lambda_{2}=1/n^{1/3}$.
$n$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.11599 0.04301 0.02163 0.01051 0.00420
$\hat{\bm{\gamma}}_{1}^{(E)}$ 1.25437 1.17177 1.13561 1.06038 0.97851
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.10449 0.03806 0.01843 0.00890 0.00361
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.10343 0.03771 0.01836 0.00888 0.00360
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.10281 0.03759 0.01842 0.00893 0.00361
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.10328 0.03796 0.01875 0.00911 0.00367
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.10634 0.03936 0.01962 0.00954 0.00383
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.14500 0.05212 0.02630 0.01272 0.00504
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.10313 0.03776 0.01838 0.00889 0.00361
$\hat{w}^{*}$ (0.17068) (0.07585) (0.04568) (0.02899) (0.01705)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.10355 0.03784 0.01839 0.00889 0.00360
$w^{*}$ (0.08453) (0.05606) (0.04310) (0.03374) (0.02475)
(b) Results when $\lambda_{1}-\lambda_{2}=1/n^{1/2}$.
$n$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.12632 0.04462 0.02185 0.01079 0.00433
$\hat{\bm{\gamma}}_{1}^{(E)}$ 1.33317 1.33026 1.31682 1.32444 1.33833
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.10748 0.03816 0.01868 0.00932 0.00363
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.10697 0.03806 0.01864 0.00929 0.00363
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.10707 0.03822 0.01872 0.00932 0.00367
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.10854 0.03889 0.01906 0.00948 0.00375
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.11343 0.04060 0.01991 0.00987 0.00393
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.15839 0.05433 0.02643 0.01296 0.00522
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.10808 0.03809 0.01865 0.00931 0.00363
$\hat{w}^{*}$ (0.17153) (0.07229) (0.04350) (0.02641) (0.01566)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.10714 0.03812 0.01866 0.00931 0.00363
$w^{*}$ (0.05375) (0.03020) (0.02054) (0.01427) (0.00895)
(c) Results when $\lambda_{1}-\lambda_{2}=1/n$.
$n$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.12627 0.04431 0.02267 0.01078 0.00442
$\hat{\bm{\gamma}}_{1}^{(E)}$ 1.44891 1.47012 1.46049 1.47838 1.48953
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.10464 0.03730 0.01857 0.00905 0.00365
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.10456 0.03733 0.01865 0.00906 0.00366
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.10508 0.03762 0.01888 0.00914 0.00371
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.10699 0.03841 0.01939 0.00934 0.00380
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.11307 0.04022 0.02044 0.00979 0.00400
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.16196 0.05414 0.02773 0.01306 0.00536
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.10721 0.03733 0.01860 0.00905 0.00365
$\hat{w}^{*}$ (0.17238) (0.07300) (0.04313) (0.02688) (0.01564)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.10461 0.03730 0.01857 0.00905 0.00365
$w^{*}$ (0.01271) (0.00440) (0.00210) (0.00102) (0.00040)
The best result up to significant digits in each setting is in bold face.
@@
We see from Table 1 that $\hat{\bm{\gamma}}_{1}(0.3)$ is the best for all cases and that $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are better than both $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}^{(T)}$. These results suggest that the optimal value of $w$, which is unknown in practice, is in $(0,1/2)$ and motivate to use the proposed estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$. As previously anticipated, $\hat{\bm{\gamma}}_{1}^{(E)}$ is not better than other estimators.
@@
We see from Table 2 that $\hat{\bm{\gamma}}_{1}(0.1)$ or $\hat{\bm{\gamma}}_{1}(0.2)$ is the best when $\eta=1/3$ and that $\hat{\bm{\gamma}}_{1}(0.1)$ or $\hat{\bm{\gamma}}_{1}^{(R)}$ or $\hat{\bm{\gamma}}_{1}(w^{*})$ or $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is the best when $\eta=1/2,1$. When $\hat{\bm{\gamma}}^{(E)}_{1}$ is consistent $(\eta<1/2)$, using small but nonzero $w$ tends to improve $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}^{(T)}$. Particularly, $\hat{\bm{\gamma}}_{1}(w^{*})$ is slightly better than $\hat{\bm{\gamma}}_{1}^{(R)}$. As for $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, it is not good as $\hat{\bm{\gamma}}_{1}(w^{*})$ in many cases, but their difference goes to 00 as $n$ increases. In contrast, $\hat{\bm{\gamma}}_{1}^{(E)}$ is quite bad especially when $\eta\geq 1/2$, which may be due to the fact from Theorem 4.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ does not necessarily satisfy (3.3). The inconsistency of $\hat{\bm{\gamma}}_{1}^{(E)}$ may affect the performance of $\hat{\bm{\gamma}}_{1}^{(T)}$, which is worse than $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, and $\hat{\bm{\gamma}}_{1}(w^{*})$.
@@
Table 3: The values of the mean squared errors (with the average of $\hat{w}^{*}$ and $w^{*}$) in the cases under large $p$, large $n$ regimes $n=\lfloor p^{0.8}\rfloor$.
(a) Results when $\lambda_{1}=p^{0.25}$, $\lambda_{2}=1$.
$p$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.10647 0.07566 0.06196 0.04910 0.03607
$\hat{\bm{\gamma}}_{1}^{(E)}$ 0.91101 0.43534 0.24993 0.16127 0.09310
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.10856 0.08709 0.07684 0.06605 0.05566
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.10488 0.08214 0.07113 0.05983 0.04856
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.10163 0.07760 0.06598 0.05436 0.04260
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.09955 0.07409 0.06199 0.05017 0.03820
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.10013 0.07276 0.06014 0.04806 0.03584
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.12513 0.08655 0.06985 0.05472 0.03944
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.10485 0.08514 0.07521 0.06475 0.05472
$\hat{w}^{*}$ (0.08214) (0.03544) (0.02728) (0.02007) (0.01245)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.10402 0.08336 0.07400 0.06406 0.05440
$w^{*}$ (0.12226) (0.07454) (0.04898) (0.03123) (0.01675)
(b) Results when $\lambda_{1}=p^{0.8}$, $\lambda_{2}=p^{0.4}$.
$p$ 20 50 100 200 500
$\hat{\bm{\gamma}}_{1}^{(T)}$ 0.21682 0.10300 0.06570 0.04191 0.02412
$\hat{\bm{\gamma}}_{1}^{(E)}$ 0.63574 0.16489 0.08581 0.04842 0.02587
$\hat{\bm{\gamma}}_{1}^{(R)}$ 0.32229 0.29513 0.30129 0.32524 0.35335
$\hat{\bm{\gamma}}_{1}(0.1)$ 0.29480 0.21403 0.16516 0.11736 0.06643
$\hat{\bm{\gamma}}_{1}(0.2)$ 0.26881 0.15606 0.10299 0.06423 0.03371
$\hat{\bm{\gamma}}_{1}(0.3)$ 0.24459 0.12276 0.07761 0.04807 0.02641
$\hat{\bm{\gamma}}_{1}(0.4)$ 0.22383 0.10712 0.06789 0.04296 0.02447
$\hat{\bm{\gamma}}_{1}(0.6)$ 0.23027 0.10690 0.06735 0.04251 0.02430
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.22046 0.10730 0.06790 0.04286 0.02451
$\hat{w}^{*}$ (0.43745) (0.43157) (0.43595) (0.44012) (0.43357)
$\hat{\bm{\gamma}}_{1}(w^{*})$ 0.21935 0.10457 0.06662 0.04241 0.02433
$w^{*}$ (0.43548) (0.43582) (0.43235) (0.42782) (0.42058)